Oven thermostats were tested by setting them to 350oF and measuring the actual
temperature of the oven. In a sample of 67 thermostats, the average temperature
was 348.2oF. If the standard deviation of the population is known to be σ = 5.1oF
do the following:
(i) Find a two-sided 95% confidence interval for the mean oven temperature.
(ii) How many thermostats must be sampled so that the 90% confidence
interval specifies the mean within ±0.8oF?
Answer #1

Solution :

(i)

Z\alpha/2 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (5.1 / \sqrt 67)

= 1.2

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

348.2 - 1.2 < \mu < 348.2 + 1.2

347.0 < \mu < 349.4

(347.0 , 349.4)

(ii)

Z\alpha/2 = 1.96

sample size = n = [Z\alpha/2* \sigma / E] 2

n = [1.96 * 5.1 / 0.8]2

n = 157

157 thermostats must be sampled .